OpenStudy (anonymous):

what is the sam as 4cm

OpenStudy (anonymous):

Which is the same as 4 cm? zero point four millimeters forty millimeters zero point zero zero four meters four hundred meters

OpenStudy (anonymous):

That is the whole question really need help to pass the exam.

OpenStudy (anonymous):

How many millimeters are in 4 centimeters? Do do this, use conversion factors: \[4cm \times \frac{10mm}{1cm} = 40mm\] Now, convert to meters: \[40mm\times\frac{1m}{1000mm} = \frac{40}{1000}m = \frac{4}{100}m = 0.04m\]

OpenStudy (anonymous):

Thanks that helped me a lot.

OpenStudy (anonymous):

I have more question that are coming later so maybe you can stick around and help me out I take FLVS for eight grade

OpenStudy (anonymous):

Is square root of seventy-one rational or irrational? Explain your reasoning